OpenStudy (babyslapmafro):

Please help me set-up the following double integral using polar coordinates (click to see).

OpenStudy (babyslapmafro):

The portion of the sphere: \[z=\sqrt{16-x^2-y^2}\] between the planes: z=1 and z=2

OpenStudy (babyslapmafro):

My problem is setting up the limits of integration.

OpenStudy (babyslapmafro):

\[\int\limits_{}^{}\int\limits_{}^{}(\sqrt{16-r^2})rdrd \theta\]

OpenStudy (anonymous):